A candy store mixes green and red jelly beans in a barrel. The barrel contains 42 pounds of mixture. If there are three times as many pounds of green jelly beans as red jelly beans, Find the number of pounds of green jelly beans and red jelly beans in the barrel.

Answer by addingup(3677)

G+R = 42
G = 3R substitute for R above:
3R+R = 42
4R = 42
R = 42/4 = 10.5 pounds of red and:
G = 3R = 3(10.5) = 31.5